Equation of motion of free falling object using lagrangian mechanics
1
Expert's answer
2020-06-08T10:28:35-0400

Let us consider the object is under free falling motion from height "H" .Mass is "m" .

At certain time the position of the object will be at "\begin{pmatrix}
 0\\
 y\\
0
\end{pmatrix}" and velocity will be "\begin{pmatrix}
 0 \\
 \dot y\\0
\end{pmatrix}" .

Now, Kinetic energy of the object will be

"K=\frac{1}{2}m\dot y^2"

And, Potential energy will be

"U=mgy"

Thus, Lagrangian is given by

"\mathscr{L}=K-U\\\implies 
\mathscr{L}=\frac{1}{2}m\dot y^2-mgy"

Therefore, from Euler Lagrange Equation of motion we get,


"\frac{\partial \mathscr{L}}{\partial y}-\frac{d}{dt}\bigg(\frac{\partial \mathscr{L}}{\partial \dot y}\bigg)=0"

Hence,

"-mg-\frac{d}{dt}(m\dot y)=0\\
\implies m\ddot y=-mg\\
\implies \boxed{ \ddot y=-g}"

Which is what exactly Newton's law predicted.